# Functional inequalities on the biased and restricted cube: an induction approach

### v2

@@ 2.3 Two-point Inequality Part
3.1 Preliminaries
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let
$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
and
$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},\quad A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
Lemma 3.1.
For any $f:A\to\mathbb{R}$, we have
Proof.
We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. Note that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
where in the second equality, we observe that
$\begin{split}\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
Lemma 3.2.
For any $f:A\to\mathbb{R}$, we have
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (25)
Proof.
Note that
$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\end{split}$
and
Thus we have
3.2 Induction part
Our goal is to prove that for every function $f:A\to\mathbb{R}$,
$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (27)
By inductive assumption, we have
Thus our goal is to show that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (30)
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (31)
Thus it suffices to show that
$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (32)
It is equivalent to show that
Lemma 3.3.
Proof.
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (34)
and
$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha)\end{split}$ (36)
Now we focus on the term
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (37)
where
$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),$
and
Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (33). ∎
3.3 Five-point Inequality Part
$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$ (40)
and
${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$ (41)
Thus to verify (33), it suffices to prove the following five-point inequality with $c=\frac{1}{2}$.
Note that
$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-\frac{a_{1}+a_{0}}{2}}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$ (43)
and
$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$ (44)
Then it suffice to prove that
$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Longleftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (45)
Note that
$a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma)$
After some simple calculations, the inequality to be verified becomes
$\begin{split}&c(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (46)
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0\end{split}$ (47)
Lemma 3.4.
$B^{2}-4AC\leq 0$.
Proof.
$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$ (48)
$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$ (49)

### v3

@@ 3.3 Two-point Inequality Part
We now show how to derive Theorem 1.7 from the $p$-biased discrete functional inequality Eq. 8 established in Theorem 1.6.
For $g:\{0,1\}^{n}\to\mathbb{R}$, recall the $p$-biased Dirichlet form
$\mathcal{E}_{p}(g,g):=\sum_{x}\mu_{p}(x)\sum_{y\sim x}\left(g(x)-g(y)\right)^{2}.$
Let $L_{p}:=D-W$ be the (external) weighted Laplacian on $A$. Then the standard edge-vertex double counting gives
$\mathcal{E}_{p}(g,g)=\sum_{\{x,y\}\subseteq A}\omega_{p}(x,y)\bigl(g(x)-g(y)\bigr)^{2}=\vec{g}^{\top}L_{p}\vec{g}.$ (38)
Indeed, grouping the two directed contributions $(x,y)$ and $(y,x)$ across each undirected edge $\{x,y\}$ yields
$\begin{split}\mu_{p}(x)(g(x)-g(y))^{2}+\mu_{p}(y)(g(x)-g(y))^{2}&=(\mu_{p}(x)+\mu_{p}(y))(g(x)-g(y))^{2}\\ &=\omega_{p}(x,y)(g(x)-g(y))^{2};\end{split}$
expanding the rightmost equality is the usual identity $\sum_{\{x,y\}}\omega_{p}(x,y)(g(x)-g(y))^{2}=\vec{g}^{\top}(D-W)\vec{g}$.
We may assume without loss of generality that $g\geq 0$. Indeed, since $||g(x)|-|g(y)||\leq|g(x)-g(y)|$, replacing $g$ by $|g|$ can only decrease the Dirichlet energy, while $\mathbb{E}_{\mu_{p}}[|g|]\geq|\mathbb{E}_{\mu_{p}}[g]|$ ensures that the right-hand side of Eq. 8 does not decrease. Therefore, Theorem 1.6 gives
$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{(\mathbb{E}_{\mu_{p}}g)^{2}}{\mu_{p}(A)}\cdot\log_{p}(\mu_{p}(A)).$ (39)
Let $\vec{\mu}=(\mu_{p}(x))_{x\in A}$, so that $\mathbb{E}_{\mu_{p}}[g]=\left\langle\vec{\mu},\vec{g}\right\rangle$. Inserting (38) into (39) gives
$\vec{g}^{\top}L_{p}\vec{g}\ \geq\rho\cdot(\vec{\mu}^{\top}\vec{g})^{2},\quad\rho:=\frac{\log_{p}\left(\mu_{p}(A)\right)}{p\mu_{p}(A)}.$
Equivalently,
$L_{p}\succeq\rho\cdot\vec{\mu}\vec{\mu}^{\top}\Rightarrow\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{1}{\rho}=\frac{p\mu_{p}(A)}{\log_{p}\big(\mu_{p}(A)\big)}.$ (40)
Since $L_{p}=D(I-P_{A})$, we have
$L_{p}^{-1}=(I-P_{A})^{-1}D^{-1}=\sum_{k\geq 0}P_{A}^{k}D^{-1}.$ (41)
Let $\vec{\nu}$ be the initial distribution $\nu(x)=\mu_{p}(x)/\mu_{p}(A)$ on $A$ and set $\vec{\phi}:=D^{-1}\vec{\mu}$, i.e. $\phi(x)=\mu_{p}(x)/d_{p}(x)$. Let $(X_{t})_{t\geq 0}$ be the Markov chain with transition $P_{A}$, killed upon exiting $A$, and let $\tau:=\min\{t\geq 1:X_{t}\notin A\}$. By the tower property and the Markov property, for every $k\geq 0$ and every $f:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$,
$(P_{A}^{k}f)(x)\ =\ \mathbb{E}_{x}\left[f(X_{k})\mathbbm{1}_{\{\tau>k\}}\right],$ (42)
where $\mathbb{E}_{x}$ is the expectation taken with respect to a $p$-biased random walk starting at a given point $x$. Applying (42) with $f=\phi$ and averaging over $\vec{\nu}$,
$\vec{\nu}^{\top}P_{A}^{k}\vec{\phi}=\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right].$
Using (41) we obtain
$\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}=\sum_{k\geq 0}\vec{\mu}^{\top}P_{A}^{k}D^{-1}\vec{\mu}=\mu_{p}(A)\sum_{k\geq 0}\vec{\nu}^{\top}P_{A}^{k}\vec{\phi}=\mu_{p}(A)\sum_{k\geq 0}\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right].$ (43)
Combining with Eq. 40 proves
$\tilde{T}(A,p)=\sum_{k\geq 0}\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right]=\frac{1}{\mu_{p}(A)}\cdot\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{p}{\log_{p}\big(\mu_{p}(A)\big)},$
where the equality case is exactly the extremal case of Theorem 1.6, namely that $A$ is an increasing subcube.
For $x$ with Hamming weight $t=|x|$, a direct calculation gives
$\phi(x)=\frac{\mu_{p}(x)}{d_{p}(x)}\ =\ \frac{1}{\,n+(n-t)\frac{p}{1-p}+t\frac{1-p}{p}\,}.$ (44)
Hence for all $x\in A$,
$\frac{\min\{p,1-p\}}{n}\ \leq\ \phi(x)\ \leq\ \frac{\max\{p,1-p\}}{n}.$ (45)
From (43) and (45) we get
$\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\geq\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\sum_{k\geq 0}\mathbb{P}_{\nu}(\tau>k)=\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\mathbb{E}_{\nu}[Y].$ (46)
Combining (40) and (46) yields
$\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\mathbb{E}_{\nu}[Y]\leq\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{p\mu_{p}(A)}{\log_{p}\big(\mu_{p}(A)\big)},$
hence
$T(A,p)=\mathbb{E}_{\nu}[Y]\leq\frac{p}{\min\{p,1-p\}}\cdot\frac{n}{\log_{p}\big(\mu_{p}(A)\big)}.$
Equality case. Equality in the final bound requires equality in each step where an inequality was used:
(a) Equality in (40) is equivalent to equality in the functional inequality (39). By the equality characterization of Theorem 1.6, this holds if and only if $A$ is an increasing subcube.
(b) Equality in (46) forces $\phi(X_{k})$ to be almost surely constant along all surviving paths. From (44) one sees that $\phi$ depends linearly on $|x|$ and is constant on AA only when $p=\tfrac{1}{2}$ (in which case $\phi\equiv 1/(2n)$). Thus the equality case occurs exactly when $p=\tfrac{1}{2}$ and $A$ is an increasing subcube.
5 Proof of mean first exit time for increasing subcubes
In this section, we verify the mean survival time of random walks from increasing subcubes (1). For $x\in A$, write $t(x):=\sum_{i=k+1}^{n}x_{i}$ for the number of free $1$’s. By permutation symmetry among the free coordinates, the conditional expectation of the exit time only depends on $t(x)$. Set
$m(t):=\mathbb{E}[Y\mid t(X_{0})=t],\ t=0,1,\dots,n-k.$
From a state with value $t$, exactly three mutually exclusive one-step outcomes can occur:
  • (exit) flip one of the first $k$ fixed $1$’s to 0 — this exits $A$;
  • (down) flip a free $1$ to 0, so $t\mapsto t-1$;
  • (up) flip a free 0 to $1$, so $t\mapsto t+1$.
Counting the three kinds of neighbors (there are $k$ exit-edges, $t$ down-edges, and nktn-k-t up-edges), we obtain
$\frac{d_{p}(x)}{\mu_{p}(x)}=\frac{k+t}{p}+\frac{n-k-t}{1-p}.$
Define
$D(t):=p(1-p)\frac{d_{p}(x)}{\mu_{p}(x)}=(1-p)(k+t)+p\bigl((n-k)-t\bigr),$
and hence the three one-step probabilities from level $t$ are
$e(t)=\frac{k(1-p)}{D(t)},\ d(t)=\frac{t(1-p)}{D(t)},\ u(t)=\frac{(n-k-t)p}{D(t)},\ e(t)+d(t)+u(t)=1.$
By the Markov property (first-step analysis), conditioning on the three one-step outcomes gives, for every $t\in\{0,\dots,n-k\}$,
$m(t)=1+d(t)\cdot m(t-1)+u(t)\cdot m(t+1),$ (47)
with d(0)=u(nk)=0d(0)=u(n-k)=0. Now let us compute the equation
$D(t)m(t)=D(t)+t(1-p)\cdot m(t-1)+(n-k-t)p\cdot m(t+1).$
Substituting $m(t)=A+Bt$, expanding both sides and matching coefficients of $t^{2},t^{1}$ and the constant term yields
$B=\frac{1-2p}{1+(1-p)k},\quad A=1+\frac{p(n-k)}{(1-p)k}(1+B).$
Under the starting law $\nu=\mu_{p}(\cdot)/\mu_{p}(A)$, the free coordinates are i.i.d. ${\rm Ber}(p)$, so $T_{0}:=t(X_{0})\sim{\rm Bin}(n-k,p)$ and $\mathbb{E}[T_{0}]=p(n-k)$. Hence
$T(C,p)=\mathbb{E}_{\nu}[Y]=\mathbb{E}[m(T_{0})]=A+B\cdot\mathbb{E}[T_{0}]=1+2p\left(\frac{n}{k}-1\right).$
This completes the proof.